OpenStudy (anonymous):

Please help! a=22, b=10, C=70 degrees. A=?

OpenStudy (lgbasallote):

sounds like cosine law :/

OpenStudy (mathmate):

Here, you have ONE angle and two adjacent sides: perfect condition for the cosine law, as Igbasallote suggested.

OpenStudy (anonymous):

use sin law \[\Large \frac{a}{\sin(A)}=\frac{b}{\sin(B)}=\frac{c}{\sin(C)}\]

OpenStudy (mathmate):

Please elaborate.

OpenStudy (anonymous):

you can find c in the above using cosine Law \[\Huge c^2=a^2+b^2-2abcos(C)\]

OpenStudy (mathmate):

Whew!

OpenStudy (anonymous):

Can you go step by step with me on this one? I don't understand :/

OpenStudy (mathmate):

Yes, in fact, we need to use the sine law to find A.

OpenStudy (mathmate):

But that would be after we have found c

OpenStudy (mathmate):

using the cosine law.

OpenStudy (anonymous):

ok using the above cosine Law \[\Large c^2=a^2+b^2-2abcos(C)\] \[\Large c^2=(22)^2+(10)^2-2(22)(10)\cos(70)\] tell me the value of c now.use your calculator.

OpenStudy (anonymous):

c^2=433.511

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take square root

OpenStudy (anonymous):

20.82

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now using the above sin law \[\Large \frac{a}{\sin(A)}=\frac{c}{\sin(C)}\] \[\Large \frac{22}{\sin(A)}=\frac{20.8}{\sin(70)}\] can you find A now ???

OpenStudy (anonymous):

20.8/sin(70)=22.13

OpenStudy (anonymous):

nopes! \[\Large \sin(A)=\frac{22*\sin(70)}{20.8}\] first solve the right side then take the sin inverse to Get A.

OpenStudy (anonymous):

Sorry, OS crashed on me. O.k. So I got .9939 and then took sin inverse which gave me 83.66

OpenStudy (anonymous):

yes that's correct !!!!!!!

OpenStudy (anonymous):

Thank you so much! I appreciate your teaching :)

OpenStudy (anonymous):

you are welcome :) any way it is really hard to get these A's everywhere (specially in Exams) :P

OpenStudy (anonymous):

Yes, it kinda is. Especially since math is my worst subject of all time! Can I ask you a question? What's the meaning of your pic? Is that your baby?